OpenStudy (anonymous):

A juggling bag is thrown straight up into the air and is caught 3s later. A: What is the initial velocity of the ball? B: Calculate the maximum height that it reaches.

OpenStudy (campbell_st):

well the initial velocity of the ball is zero as you are juggling bags... and it has no max height... as you again are using bags

OpenStudy (mrnood):

@Campbell_st That is not a helpful answer @fjapusen I suggest you ignore that post

OpenStudy (mrnood):

You need the equations for constant acceleration. In tis case you know: a = g = 9.81m/s^2 t= 1.5s (because the bag will take the same time to reach peak as to fall back to catching point) v (final velocity) = 0 because the bag stops at the top the equation to solve this is v= u +at (solve for u using the above values) the other equation required is s=ut +0.5 at^2 solve this for s (after you have the value for u from the first part)